### v3

# From Words to Collisions: LLM-Guided Evaluation and Adversarial Generation of Safety-Critical Driving Scenarios

Abstract
Ensuring the safety of autonomous vehicles requires virtual scenario-based testing, which depends on the robust evaluation and generation of safety-critical scenarios. So far, researchers have used scenario-based testing frameworks that rely heavily on handcrafted scenarios as safety metrics. To reduce the effort of human interpretation and overcome the limited scalability of these approaches, we combine Large Language Models (LLMs) with structured scenario parsing and prompt engineering to automatically evaluate and generate safety-critical driving scenarios. We introduce Cartesian and Ego-centric prompt strategies for scenario evaluation, and an adversarial generation module that modifies trajectories of risk-inducing vehicles (ego-attackers) to create critical scenarios. We validate our approach using a 2D simulation framework and multiple pre-trained LLMs. The results show that the evaluation module effectively detects collision scenarios and infers scenario safety. Meanwhile, the new generation module identifies high-risk agents and synthesizes realistic, safety-critical scenarios. We conclude that an LLM equipped with domain-informed prompting techniques can effectively evaluate and generate safety-critical driving scenarios, reducing dependence on handcrafted metrics. We release our open-source code and scenarios at: https://github.com/TUM-AVS/From-Words-to-Collisions.
Index Terms:
Autonomous Driving, Large Language Models, Scenario-based Test, Safety-critical Scenario Evaluation, Safety-critical Scenario Generation
I Introduction
The development and deployment of autonomous vehicles have progressed rapidly, reducing human intervention within specific Operational Design Domains (ODDs) step by step. Companies like Waymo have deployed fully autonomous SAE Level 4 [1] robotaxi services in defined ODDs, demonstrating the potential of driverless technology in urban environments. This advancement is primarily driven by developing and validating highly reliable Automated Driving Systems (ADS). Traditionally, validation has relied on real-world testing, including on-road trials [2]. However, real-world testing cannot fully capture the diversity of driving situations and edge cases. To address this, researchers and industry have increasingly employed virtual scenario-based testing, enabling cost-effective simulation of realistic conditions [3].
Figure 1: Our paper presents a new LLM-based framework with two modules: Module I: LLM-based evaluation of safety-critical driving scenarios. Module II: LLM-based generation of safety-critical driving scenarios.
Scenario-based testing needs to evaluate the behavior of ADS in safety-critical contexts. Various metrics have been proposed to assess the safety-criticality of driving scenarios [4], including proximity-based indicators like Time-to-Collision (TTC) and Minimum Distance to Collision (MDC), categorized into temporal and non-temporal types. Based on these metrics, test engineers can evaluate the safety-criticality of a scenario.
Since the release of GPT-3 [5], applications using Large Language Models (LLMs) have surged. As large-scale pre-trained models on general-purpose datasets [6], LLMs combine strong generalization with adaptability, allowing direct application to domain-specific tasks via prompting or fine-tuning [7]. Recently, researchers have explored their use in autonomous driving tasks, including perception [8], motion planning [9], and vehicle control [10]. This raises a key question: Given their good generalization and adaptability, can LLMs be used to evaluate and generate safety-critical driving scenarios, as illustrated in Fig. 1?
II Related Work
Evaluating and generating safety-critical scenarios are key to scenario-based testing in autonomous driving. This section reviews both traditional and LLMs-based approaches.
II-A Safety-Critical Driving Scenario Evaluation for ADS
Several recent works have explored evaluating safety-critical driving scenarios using different methods. In [11], safety-critical scenarios were defined as collisions or near-collisions with the ego vehicle. In [12], a set of safety metrics was employed in scenario evaluation, including TTC, braking behavior, required deceleration, and traffic-related indicators. While these safety metrics can be computed from scenario data, analyzing them to assess safety requires human effort and domain knowledge.
LLMs as evaluator: Recent studies have explored LLMs for driving scenario evaluation, focusing on realism, behavior, and consistency. In [13], they prompted LLMs with contextual data, e.g., road layout, weather, vehicle behavior, to assess scenario realism using the DeepScenario dataset [14]. [15] proposes a framework incorporating high-level evaluation criteria, including performance, safety, and comfort, alongside scenario data from the CARLA [16] to evaluate driving style and performance. However, in [15] they focus on assessing final driving style and driving performance level based on aggregated input, rather than directly assessing individual metrics like safety. OmniTester [17] extended text-only evaluation to multimodal analysis by generating captions from visualized scenarios using a vision-language model and comparing them with original descriptions from LLM.
II-B Safety-Critical Driving Scenario Generation for ADS
Many studies have explored the generation of safety-critical scenarios to test ADS. According to the survey [18], the main approaches are data-driven, adversarial, and knowledge-based. Data-driven methods use real-world datasets to generate realistic scenarios and avoid overfitting, leveraging models like Bayesian networks [19] and deep generative models [20]. Adversarial approaches aim to expose weaknesses in ADSs by synthesizing rare but high-risk situations using techniques such as differentiable renderers [21] or reinforcement learning [22]. Knowledge-based methods integrate domain expertise, e.g., traffic rules, through manual rules [23] or hybrid approaches combining rule-based priors with adversarial policy learning [24] for enhanced controllability.
LLMs as generator: Recently, LLMs have emerged as powerful tools for generating simulation scenarios directly from natural language. Depending on the abstraction level and application, the underlying simulators fall into two categories: microscopic, such as CARLA [16], focusing on ego vehicle behavior, and macroscopic, like SUMO [25], enabling large-scale traffic simulations. At the microscopic level, ChatScene [26] and TTSG [27] used LLMs to generate safety-critical scenarios in CARLA from textual descriptions. ChatScene further utilized these scenarios for training and evaluating vehicle control algorithms, while TTSG applied them to multi-agent planning. For macroscopic scenario generation, ChatSUMO [28] and OmniTester [17] leveraged LLMs to produce realistic urban traffic simulations from text, showcasing the scalability of language-driven scenario generation.
II-C Critical Summary
To the best of our knowledge, the existing literature is limited by at least one of the following aspects:
  1. 1.
    The safety-criticality of scenarios is typically evaluated by test engineers using predefined metrics. For example, ChatScene [26] and TTSG [27] computed metrics such as TTC and MDC, but then a human had to interpret the generated scenarios.
  2. 2.
    No examples of LLMs to evaluate safety-critical scenarios. The authors of [13] and [17] analyzed the scenario realism and semantic consistency, while [16] evaluated the overall driving style and performance. However, no one has performed any safety-criticality assessment so far.
  3. 3.
    Existing LLM-based scenario generators relied on textual descriptions provided by experts. Integrating adversarial methods into LLMs to generate safety-critical scenarios remains unexplored.
II-D Contributions
To address the previous limitations, the key contributions of this paper are the following:
  • We present a novel framework that integrates LLMs as evaluation modules to assess the safety-criticality of structured driving scenarios.
  • We compare different prompt formulations to analyze safety-critical scenarios and focus on two use cases: collision evaluation and safety inference.
  • We propose a novel LLM-based module that transforms safe scenarios into safety-critical ones by adversarially modifying the trajectories of selected agents.
  • We present empirical evaluations through ablation studies on different prompt formulations and multiple state-of-the-art LLMs, tested on 200 randomly selected scenarios from a dataset of 6000 simulations.
Figure 2: Overview of the proposed framework based on LLM to evaluate and generate safety-critical scenarios.
III Methodology
This paper proposes a novel LLM-based framework (Fig. 2) for evaluating and generating safety-critical driving scenarios. Primarily, we focus on scenarios from a 2D simulator like CommonRoad[29], which supports scenario-based tests for the motion planning algorithms.
Module I-Evaluation (top parts of Fig. 12): In the LLM-based scenario evaluation module, a key challenge is that LLMs require natural language input, making direct processing of structured scenario data difficult. To address this, we introduce a parser converting structured data into natural language descriptions, serving as contextual input for the LLM. To enhance domain-specific adaptability for scenario evaluation and safety inference, we propose and compare Cartesian and Ego coordinate prompt templates.
Module II-Generation (bottom parts of Fig. 12): We propose an LLM-based generation module that transforms safe scenarios into safety-critical ones. Our framework identifies potential ego-attackers by analyzing the agents’ motion and safety-critical metrics (TTC and MDC), provided through prompts. Then, our model modifies the ego-attackers’ trajectories adversarially to induce collisions with the Ego vehicle (EV).
III-A Structured Driving Scenario Data
This section describes the driving scenario data used as input by our framework ("Input" and "Structured Scenario Data" blocks in Fig. 2). We employ a 2D simulator providing a standardized scenario representation, which contains information on the road map, road network, traffic signs, traffic participants, and motion planning task. This structure supports the testing and benchmarking motion planning algorithms across diverse driving scenarios.
In our scenario representation, the road network uses the lanelet map [30], which is aligned with OpenStreetMap (OSM) standards, and includes detailed lane boundaries, connectivity, and traffic flow topology. The driving scenario incorporates traffic infrastructure elements (e.g., signs, traffic lights) and various traffic participants, such as cars, buses, and motorcycles, which are classified as static or dynamic obstacles. All entities are annotated with geometric and kinematic attributes to simulate real-world interactions accurately. The scenario data also includes the inputs for a motion planning problem, which specifies the EV ’s initial state, goal states, and the planning task (Fig. 3).
Figure 3: Example of an overtaking scenario in the 2D simulator with the motion planner used in this paper1.
footnotetext: We use the CommonRoad simulator [29] and the Frenetix motion planner [31], but other 2D simulators and motion planner could be equivalently used.
To generate a dataset of scenarios for our framework, we perform a set of multi-vehicle open-loop simulations, in which the EV’s motion planner tries to generate collision-free trajectories to reach a planning goal. We collect these simulation scenarios in a structured dataset, which will be used as input by our framework. We define our structured scenario data as:
$S=\{\mathcal{L},\mathcal{T},\mathcal{O},\mathcal{E}\}$ (1)
where $\mathcal{L}$ represents the lanelets, $\mathcal{T}$ denotes the traffic signs, $\mathcal{O}$ is the set of obstacles, and $\mathcal{E}$ represents the EV state.
III-B LLM-based Safety-Critical Scenario Evaluation Module
III-B1 Parser
To bridge the gap between the structured scenario dataset $\mathcal{S}$ and the natural language input required by the LLM, we introduce a parser that transforms structured scenario data into textual descriptions (Fig. 2). Specifically, we define a parsing function $\mathcal{F}$ that maps scenario data $\mathcal{S}$ to a context $\mathcal{C}$, such that $\mathcal{C}=\mathcal{F}(\mathcal{S})$.
This parsing function extracts key elements from $\mathcal{S}$, including lanelets information $\mathcal{L}$, obstacles states $\mathcal{O}$, and EV states $\mathcal{E}$, excluding traffic signs $\mathcal{T}$, which are not required for evaluation. Obstacles and EV data are represented in global Cartesian coordinates, including position, velocity, orientation, and acceleration.
We design two specialized parsers: one using a map-based representation in Cartesian coordinates, named $\mathcal{F}_{\text{cart}}$, and the other using an ego-centric description, named $\mathcal{F}_{\text{ego}}$. $\mathcal{F}_{\text{cart}}$ describes the absolute states of all agents in Cartesian coordinates, and its output is illustrated in Fig. 4(a) In contrast, the Ego-centric parser $\mathcal{F}_{\text{ego}}$ describes the scenario from the EV’s perspective, encoding relative spatial relations (e.g., front-rear, left-right) and dynamics (e.g., position, velocity, acceleration) using the longitudinal and lateral coordinates $\{s,d\}$. It also generates relative motion descriptions for each obstacle, such as “$m_{s}$: The obstacle is approaching the EV longitudinally from the front”, “$m_{d}$: The obstacle is moving laterally towards the EV from the left”. The output of $\mathcal{F}_{\text{ego}}$ is shown in Fig. 4(b). More details about these motion descriptions can be found in our open-source code repository. The outputs of $\mathcal{F}_{\text{cart}}$ and $\mathcal{F}_{\text{ego}}$ provide the context for the following Cartesian and Ego prompts, described in the next subsection.
Figure 4: Output from the Cartesian-coordinate parser (a) and Ego-centric parser (b). The blue part describes the ego vehicle’s state, the orange parts are the lanelet’s identification numbers, and the green part describes the obstacles’ states.
III-B2 Prompt Formulation
Since LLMs are pre-trained on vast amounts of unlabeled data, techniques like prompt engineering or fine-tuning [7] are commonly used to adapt the LLMs to downstream tasks. Given the strong reasoning capabilities of state-of-the-art models such as GPT-4o [32], prompt-based adaptation is often favored over fine-tuning. In this prompt approach, a system message defines the model’s behavior and domain, while a user message specifies task instructions, enabling adaptation without changing the model’s weights. Specifically, we use prompting to adapt LLMs to evaluate safety-critical scenarios.
We propose a structured prompting template integrating advanced techniques with Contextual Prompting, Chain-of-Thought reasoning, and In-Context Learning. Within this template, we define two prompt types based on the contextual input: the Cartesian coordinate prompt ($\textit{Prompt}_{\mathrm{cart}}$) and the Ego coordinate prompt ($\textit{Prompt}_{\mathrm{ego}}$). In both cases, a system message assigns the LLM the role of a collision evaluation expert, enabling reasoning over safety-critical metrics such as TTC and MDC. The key difference lies in the user message, which considers the scenario information based on Cartesian or Ego coordinates (Fig. 4). In Section IV, we compare the performance of the two prompts. The following paragraphs detail the techniques that we employ in our prompt design.
Contextual Prompting (CP): The system message defines key metrics, such as TTC and MDC, and quantifies safety-criticality by assigning risk scores based on threshold values for these metrics. By embedding this domain-specific knowledge, CP enhances the model’s understanding of evaluation criteria, improving the reliability of its responses. Furthermore, the framework remains flexible, allowing additional metrics to be easily incorporated.
Chain-of-Thought (CoT): CoT prompting explicitly guides the LLM to reason step-by-step rather than providing immediate answers. In our framework, the user message requests intermediate reasoning steps such as assessing obstacles’ trajectories, identifying critical obstacles, and evaluating the scenario’s safety-criticality.
In-Context Learning (ICL): This prompt incorporates examples of collision and risk assessments within the system message. These demonstrations guide the LLM’s reasoning process, define the task format and evaluation logic.
By integrating the CP, CoT, and ICL techniques within $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$ and $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$, we adapt a general-purpose LLM to a specialized evaluator of driving scenarios. In Section IV, ablation studies will assess how these techniques impact the model’s performance.
III-C LLM-Based Safety-Critical Scenario Generation Module
We extend our framework with an adversarial scenario generation module that transforms safe driving scenarios into safety-critical ones. Our method identifies potential ego-attackers and adversarially changes their trajectories, based on agent-level risk scores.
III-C1 Parser
To identify ego-attackers, we directly integrate the values of the TTC and MDC safety-critical metrics into the evaluation prompt. For this purpose, we design a safety-critical metrics parser that converts structured metric data into natural language context for the LLM, as shown in Fig. 5(a). Once an ego-attacker is identified, we use a trajectory generation parser to convert its recorded behavior into textual format, illustrated in Fig. 5(b).
III-C2 Prompt Formulations
In the generation of safety-critical scenarios, we combine two prompts: a safety-critical metrics prompt, to assess the obstacles’ risk scores, and a trajectory generation prompt, to synthesize new adversarial trajectories for the ego attackers.
The safety-critical metrics prompt follows the structure of $\textit{Prompt}_{\mathrm{cart}}$ and $\textit{Prompt}_{\mathrm{ego}}$, with the LLM acting as a domain expert in collision evaluation. It incorporates CoT reasoning and ICL examples to guide the assessment of overall risk scores for each obstacle, based on thresholds for each safety-critical metric. The prompt also includes motion descriptions, as shown in Fig. 5(a), focusing on obstacles approaching from the front or sides, since these directions are more relevant for potential collision scenarios crucial to evaluating planner performance. This enhanced structure enables the LLM to systematically reason about ego-attackers using both quantitative metrics and spatial context.
For trajectory generation, we adopt a different prompting strategy as the LLM transitions from evaluator to generator. The system message instructs the model to synthesize an adversarial trajectory that increases the collision likelihood with the EV, given the scenario context. To enhance the generation performance, we apply Self-Consistency (SC) by prompting the model to produce multiple adversarial trajectory candidates for the same ego-attacker. Each candidate is evaluated using the safety-critical metrics prompt, and the trajectory with the lowest (i.e., most critical) risk score for the ego-attacker is selected.
Figure 5: Output from safety-critical metrics parser (a) and trajectory generation parser (b). The blue part describes the ego state, the orange parts are the lanelet identification numbers, and the green parts are the obstacles’ states.
III-D Structured Output Format
In our LLM-based scenario evaluation and generation framework, we enforce structured output formats using predefined schemes, which ensure consistency and reduce hallucinated outputs [33]. These formats define fixed keys aligned with specific use cases, including collision evaluation, safety-critical scenario inference, agent-level risk assessment (e.g., risk scores), and trajectory generation parameters (e.g., position, velocity). The structure remains flexible and can be adapted to various prompt designs and task requirements.
IV Results & Discussion
IV-A Experimental Setup
We conduct our experiments using the 2D simulator CommonRoad[29], with the Frenetix [31] open-source motion planner. This simulation setup enables us to collect diverse driving scenario datasets to be used as input for our framework. In our experiments, we compare the latest LLMs Gemini-1.5Pro [34], GPT4o, and DeepSeek-V3 [35], which are the top powerful LLMs in the open leaderboard, using their APIs.
All experiments are performed on a Dell Alienware R15 equipped with an Intel i7-13700KF CPU, a NVIDIA RTX 4090 GPU, and a 64 GB RAM. We simulate 6000 real recorded driving scenarios across different countries’ maps, road categories, obstacle classes, and number of traffic participants. The Frenetix motion planner successfully generates collision-free trajectories in over 4700 scenarios, allowing the EV to reach its goal state. In approximately 1000 scenarios, the EV collides with a traffic participant. The remaining scenarios involve other failures or exceptional conditions, such as unsolvable cases. We use the collision cases to evaluate our scenario evaluator, treating them as ground truth for the collision evaluation use case. For the safety inference use case, we further identify safety-critical obstacles by computing TTC and MDC with risk score as ground truth. Lastly, for the scenario generation use case, we assess our generator by transforming safe scenarios into safety-critical ones through adversarial trajectory modifications.
Figure 6: Confusion matrices for collision evaluation using $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$ and $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$.
IV-B Collision Evaluation in Driving Scenarios
This use case evaluates if collisions occur in driving scenarios with EV. The inputs of our LLM model include a comprehensive map with geometric lanelet information, dynamics status of the EV, and of all traffic participants throughout the simulation’s duration. Our model outputs a boolean flag to indicate if a collision occurs for the EV in the given scenario. We randomly select 100 non-collision and 100 collision scenarios for evaluation from the 6000 total scenarios.
To test our LLM-based collision evaluator, we employ the two prompt types $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$ and $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$ [Section III-B2] with collision evaluation output format [Section  III-D], using the Gemini-1.5Pro model. The results are summarized in Fig. 6, which presents the confusion matrices for each prompt setting. $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$ fails to detect most collisions, identifying only 12% of collision scenarios correctly and resulting in an 88% false negative rate. This indicates that Cartesian-based input is insufficient for accurate scenario risk assessment. In contrast, $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$ achieves 96% accuracy on collisions and 80% on non-collisions, demonstrating the effectiveness of ego-centric inputs for safety-critical evaluation from the EV’s perspective.
IV-B1 Ablation Study
The performance of our LLM-based evaluator depends on prompt design and advanced techniques with Contextual Prompting (CP), In-Context Learning (ICL), and Chain-of-Thought (CoT) reasoning, as described in Section III-B2. To evaluate their impact, we conduct an ablation study comparing four configurations for both $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$ and $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$: Base (no techniques), CP (CP only), CoT (CP + CoT), and All (CP + CoT + ICL).
TABLE I: Ablation study on collision evaluation.
Prompt Method TP FP FN TN Precision Recall F1 Score
Cart Base 1 0 99 100 1.00 0.01 0.02
Cart CP 2 0 98 100 1.00 0.02 0.04
Cart CoT 6 0 94 100 1.00 0.06 0.11
Cart All 12 0 88 100 1.00 0.12 0.21
Ego Base 54 0 46 100 1.00 0.54 0.70
Ego CP 66 8 34 92 0.89 0.66 0.79
Ego CoT 75 8 25 92 0.90 0.75 0.84
Ego All 96 20 4 80 0.83 0.96 0.87
Cart and Ego denote $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$ and $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$, respectively. TP (True Positive): collision correctly identified; FP (False Positive): non-collision incorrectly identified as collision; FN (False Negative): collision missed; TN (True Negative): non-collision correctly identified.
TABLE II: LLMs comparison for collision evaluation.
Prompt LLM TP FP FN TN Precision Recall F1 Score
Cart Gemini-1.5Pro 12 0 88 100 1.00 0.12 0.21
Cart GPT-4o 1 0 99 104 1.00 0.01 0.02
Cart DeepSeek-V3 2 2 98 102 0.50 0.02 0.04
Ego Gemini-1.5Pro 96 20 4 80 0.83 0.96 0.87
Ego GPT-4o 95 20 5 84 0.83 0.95 0.88
Ego DeepSeek-V3 97 19 3 85 0.84 0.97 0.90
As shown in Table I, $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$ performs poorly across all settings, with a maximum F1 score of 0.21, underscoring its limitations in collision reasoning. In contrast, $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$ shows clear gains with each added prompting technique, improving from an F1 score of 0.70 (Base) to 0.87 (All). The All configuration achieves the highest recall (0.96), demonstrating strong detection of true collisions. While precision drops slightly from 0.89 to 0.83, this reflects a typical trade-off favoring recall in safety-critical tasks. Overall, the results confirm that CP, CoT, and ICL significantly enhance the evaluator’s performance.
IV-B2 Comparison Across Different LLMs
We compare the reasoning performance of Gemini-1.5Pro, GPT-4o, and DeepSeek-V3 using the same $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$ and $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$ with All configuration applied. As shown in Table II, all models perform significantly better with $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$, reinforcing the findings in Section IV-B1. Gemini-1.5Pro achieves the highest F1 score (0.21) with $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$, likely due to its larger context window (up to 200k tokens), which helps interpret verbose Cartesian inputs. With ego-centric prompts, DeepSeek-V3 with an F1 score of 0.90 slightly outperforms the others, followed by GPT-4o (0.88) and Gemini-1.5Pro (0.87), suggesting a slight edge in reasoning ability for DeepSeek-V3.
This evaluation approach can be readily extended to general safety-criticality assessments, such as classifying scenarios as safe or high-risk, as it only requires structured scenario data. A generalized example is available in our GitHub repository. Since traditional metric-based methods are used to generate the ground truth, we do not perform direct comparisons against them.
IV-C Safety Inference in Driving Scenarios
[…]
Figure 7: Confusion matrices for safety evaluation, using $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$ and $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$.
IV-C1 Ablation Study
We conduct an ablation study using four configurations Base, CP, CoT, and All applied to both $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$ and $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$ to assess their impact on safety inference. Table III shows that $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$ consistently outperforms $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$. Notably, $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$ achieves solid performance in the safety inference task, with the All configuration reaching an F1 score of 0.62. This result suggests that Cartesian inputs can be effective when the task is limited to nearby agents and local context.
TABLE III: Ablation study on safety inference.
Prompt Method TP FP FN TN Precision Recall F1 Score
Cart Base 4 18 96 82 0.18 0.04 0.07
Cart CP 11 6 89 94 0.65 0.11 0.19
Cart CoT 42 8 58 92 0.84 0.42 0.56
Cart All 45 1 55 99 0.98 0.45 0.62
Ego Base 64 23 36 77 0.74 0.64 0.69
Ego CP 65 24 35 76 0.72 0.65 0.68
Ego CoT 73 16 27 84 0.82 0.73 0.77
Ego All 84 18 16 82 0.82 0.84 0.83
TABLE IV: LLMs comparison for safety inference.
Prompt LLM TP FP FN TN Precision Recall F1 Score
Cart Gemini-1.5Pro 45 1 55 99 0.98 0.45 0.62
Cart GPT-4o 22 9 78 91 0.71 0.22 0.34
Cart DeepSeek-V3 37 2 63 98 0.95 0.37 0.53
Ego Gemini-1.5Pro 84 18 16 82 0.82 0.84 0.83
Ego GPT-4o 85 24 15 76 0.78 0.85 0.81
Ego DeepSeek-V3 89 16 11 84 0.85 0.89 0.87
However, the $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$ yields strong and balanced results across all settings, with the All configuration achieving the highest F1 score (0.83) and recall (0.84). These findings reinforce that ego-centric prompts, especially when enhanced by CP, ICL, and CoT techniques, are most effective for safety inference.
IV-C2 Comparison Across Different LLMs
To assess the reasoning performance of different LLMs in the safety inference task, we apply the same $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$ and $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$ using the All prompting configuration. The results are summarized in Table IV. All models perform significantly better with $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$, reaffirming the value of ego-centric representations. DeepSeek-V3 achieves the highest F1 score (0.87), followed by Gemini-1.5Pro (0.83) and GPT-4o (0.81). On Cartesian input, Gemini-1.5Pro leads with an F1 score of 0.62, outperforming GPT-4o (0.34) and DeepSeek-V3 (0.53), likely due to its larger context window, which helps interpret verbose global inputs. In contrast, DeepSeek-V3 slightly outperforms the others in terms of ego-centric input. This suggests that model reasoning performance plays a greater role when context length becomes less critical.
Overall, our safety inference framework operates independently during simulation and can serve as a supplementary module to support motion planners in online risk assessment. While the LLM-based evaluator shows strong reasoning capabilities, its current response time, averaging several seconds per inference, remains unsuitable for real-time deployment. Future work may explore fine-tuning smaller, task-specific models to enable faster, safer inference. Additional runtime details and performance metrics are provided in our GitHub repository.
IV-D Generation of Safety-Critical Scenarios
[…]
Figure 8: Risk score of the ego-attacker vehicle’s trajectory, in the original safe scenario and in the generated safety-critical one. A risk score 5 indicates very low risk, while 0 denotes a collision.
Figure 9: A case study of safety-critical scenario generation on CommonRoad[29] BEL Antwerp-1_14_T-1. The ego vehicle is in blue, and the ego-attacker vehicle is in red. The top row depicts the original safe scenario, while the bottom row shows the corresponding safety-critical scenario generated by our model.
To assess the effectiveness of our framework, we conduct a case study using a scenario from the previous 4,700 safe scenarios from Sec IV-A. As shown in Fig. 8 and 9, the original scenario is safe, and the ego-attacker’s trajectory has a safe risk level. Conversely, in the generated scenario, the ego-attacker’s modified trajectory has a decreasing risk score over time (Fig. 8), ultimately leading to a collision. Fig. 9 further visualizes this evolution. The top row shows the original non-critical scenario, where the blue EV and the red ego-attacker navigate safely. The bottom row illustrates the modified scenario, where a collision is triggered. These results validate the ability of our framework to generate targeted, safety-critical scenarios based on structured LLM-driven risk inference.
Since we focus on identifying the ego-attacker and generating safety-critical scenarios adversarially, our framework can provide greater efficiency and controllability compared to purely text-based scenario generation approaches integrating LLMs.
V Conclusion & Outlook
This paper introduced a novel LLM-based framework for evaluating and generating safety-critical driving scenarios from structured data. We proposed two prompt templates $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$ and $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$ supported by parsers, and advanced prompting techniques including Contextual Prompting, Chain-of-Thought reasoning, and In-Context Learning. We also developed an LLM-guided scenario generation module to generate safety-critical scenarios by adversarially modifying ego-attackers’ trajectories to create the safety-critical cases.
The framework was evaluated in three use cases supported by structured outputs: collision scenario evaluation, safety inference, and safety-critical scenario generation. Across all LLMs, $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$ consistently outperformed $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$, for the Gemini-1.5Pro, achieving F1 scores of 0.87 vs. 0.21 in collision evaluation, and 0.83 vs. 0.62 in safety inference, highlighting the advantage of ego-centric representations for reasoning about risk. In generation, the framework successfully minimized the ego-attacker’s risk score to synthesize safety-critical outcomes. Future work includes extending the framework to 3D simulators such as CARLA, and enhancing the generation module to produce more diverse and challenging safety-critical scenarios.

### v4

@@
 OmniTester [16] extended text-only evaluation to multimodal analysis by generating captions from visualized scenarios using a vision-language model and comparing them with original descriptions from LLM.
+Notably, a recent survey [17] summarizes existing LLM-based scenario analysis methods.
+However, none of the papers explicitly assess safety criticality in driving scenarios by using LLMs.
 II-B Safety-Critical Driving Scenario Generation for ADS
@@
 $\mathcal{F}_{\text{cart}}$ describes the absolute states of all agents in Cartesian coordinates, and its output is illustrated in Fig.
-4(a) In contrast, the Ego-centric parser $\mathcal{F}_{\text{ego}}$ describes the scenario from the EV’s perspective, encoding relative spatial relations (e.g., front-rear, left-right) and dynamics (e.g., position, velocity, acceleration) using the longitudinal and lateral coordinates $\{s,d\}$.
+4 (a) In contrast, the Ego-centric parser $\mathcal{F}_{\text{ego}}$ describes the scenario from the EV’s perspective, encoding relative spatial relations (e.g., front-rear, left-right) and dynamics (e.g., position, velocity, acceleration) using the longitudinal and lateral coordinates $\{s,d\}$.
 It also generates relative motion descriptions for each obstacle, such as “$m_{s}$: The obstacle is approaching the EV longitudinally from the front”, “$m_{d}$: The obstacle is moving laterally towards the EV from the left”.
@@
 It incorporates CoT reasoning and ICL examples to guide the assessment of overall risk scores for each obstacle, based on thresholds for each safety-critical metric.
-The prompt also includes motion descriptions, as shown in Fig.
-5(a), focusing on obstacles approaching from the front or sides, since these directions are more relevant for potential collision scenarios crucial to evaluating planner performance.
-This enhanced structure enables the LLM to systematically reason about ego-attackers using both quantitative metrics and spatial context.
+Our prompt also includes motion descriptions, as shown in Fig.
+5(a), focusing on obstacles approaching from the front or sides, since these directions are relevant to evaluate the planner’s performance.
+This structure enables the LLM to reason about ego-attackers using both quantitative metrics and spatial context.
 For trajectory generation, we adopt a different prompting strategy as the LLM transitions from evaluator to generator.
@@
 To enhance the generation performance, we apply Self-Consistency (SC) by prompting the model to produce multiple adversarial trajectory candidates for the same ego-attacker.
-Each candidate is evaluated using the safety-critical metrics prompt, and the trajectory with the lowest (i.e., most critical) risk score for the ego-attacker is selected.
+Each candidate is scored using the safety-critical metrics prompt, and the one with a low-risk score for the ego-attacker is selected.
 Figure 5: Output from safety-critical metrics parser (a) and trajectory generation parser (b).
@@
 III-D Structured Output Format
-In our LLM-based scenario evaluation and generation framework, we enforce structured output formats using predefined schemes, which ensure consistency and reduce hallucinated outputs [33].
-These formats define fixed keys aligned with specific use cases, including collision evaluation, safety-critical scenario inference, agent-level risk assessment (e.g., risk scores), and trajectory generation parameters (e.g., position, velocity).
+In our LLM-based scenario evaluation and generation framework, we enforce structured output formats using predefined schemes to ensure consistency and reduce hallucination [33].
+These formats define fixed keys aligned with specific use cases, including collision evaluation, safety-critical scenario inference, agent-level risk assessment (e.g., risk scores), and trajectory generation parameters (e.g., position).
 The structure remains flexible and can be adapted to various prompt designs and task requirements.
@@
 IV-B Collision Evaluation in Driving Scenarios
-This use case evaluates if collisions occur in driving scenarios with EV.
-The inputs of our LLM model include a comprehensive map with geometric lanelet information, dynamics status of the EV, and of all traffic participants throughout the simulation’s duration.
+This use case evaluates if collisions occur with the EV in driving scenarios.
+The inputs of our LLM include a map with geometric lanelet information, dynamic states of the EV, and of all traffic participants throughout the simulation’s duration.
 Our model outputs a boolean flag to indicate if a collision occurs for the EV in the given scenario.
@@
 IV-B1 Ablation Study
-The performance of our LLM-based evaluator depends on prompt design and advanced techniques with Contextual Prompting (CP), In-Context Learning (ICL), and Chain-of-Thought (CoT) reasoning, as described in Section III-B2.
-To evaluate their impact, we conduct an ablation study comparing four configurations for both $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$ and $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$: Base (no techniques), CP (CP only), CoT (CP + CoT), and All (CP + CoT + ICL).
+The performance of our LLM-based evaluator depends on prompt design and advanced techniques with CP, ICL, and CoT reasoning, as described in Section III-B2.
+To evaluate their impact, an ablation study is conducted for both $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$ and $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$: Base (no techniques), CP (CP only), CoT (CP + CoT), and All (CP + CoT + ICL).
 TABLE I: Ablation study on collision evaluation.
@@
 We compare the reasoning performance of Gemini-1.5Pro, GPT-4o, and DeepSeek-V3 using the same $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$ and $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$ with All configuration applied.
-As shown in Table II, all models perform significantly better with $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$, reinforcing the findings in Section IV-B1.
+As shown in Table II, all models perform significantly better with $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$, confirming the findings in Section IV-B1.
 Gemini-1.5Pro achieves the highest F1 score (0.21) with $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$, likely due to its larger context window (up to 200k tokens), which helps interpret verbose Cartesian inputs.
-With ego-centric prompts, DeepSeek-V3 with an F1 score of 0.90 slightly outperforms the others, followed by GPT-4o (0.88) and Gemini-1.5Pro (0.87), suggesting a slight edge in reasoning ability for DeepSeek-V3.
+With ego-centric prompts, DeepSeek-V3 slightly outperforms the others (F1 score 0.90), followed by GPT-4o (0.88) and Gemini-1.5Pro (0.87).
 This evaluation approach can be readily extended to general safety-criticality assessments, such as classifying scenarios as safe or high-risk, as it only requires structured scenario data.
@@
 Additional runtime details and performance metrics are provided in our GitHub repository.
+0022224444666688881010101012121212141414140011112222333344445555Refer to captionTimestepRisk Score (0-5)Original safe scenarioGenerated safety-critical scenario
+Figure 8: Risk score of the ego-attacker vehicle’s trajectory, in the original safe scenario and in the generated safety-critical one.
+A risk score of 5 indicates very low risk, while 0 denotes a collision.
 IV-D Generation of Safety-Critical Scenarios
 […]
-Figure 8: Risk score of the ego-attacker vehicle’s trajectory, in the original safe scenario and in the generated safety-critical one.
-A risk score 5 indicates very low risk, while 0 denotes a collision.
 Figure 9: A case study of safety-critical scenario generation on CommonRoad[29] BEL Antwerp-1_14_T-1.
@@
 In generation, the framework successfully minimized the ego-attacker’s risk score to synthesize safety-critical outcomes.
-Future work includes extending the framework to 3D simulators such as CARLA, and enhancing the generation module to produce more diverse and challenging safety-critical scenarios.
+Future work includes extending the framework to 3D simulators such as CARLA, enhancing the generation module to produce more diverse safety-critical scenarios, and using the generated scenarios to further evaluate motion planning algorithms [36].